OpenStudy (anonymous):

How to solve this differential equation: (1+x)*2 dy/dx+y*2=1

OpenStudy (amistre64):

looks separable to me ....

OpenStudy (amistre64):

i take it you used * instead of ^, but meant ^

OpenStudy (anonymous):

Yes.

OpenStudy (amistre64):

do you know how to do separable stuff?

OpenStudy (anonymous):

Yes, but I just want to check my steps whether its correct.

OpenStudy (amistre64):

show me what you got, ill see if its proper

OpenStudy (amistre64):

\[f(x)\frac{dy}{dx}+g(y)=c\] \[f(x)\frac{dy}{dx}=c-g(y)\] \[f(x)dy=[c-g(y)]dx\] \[\frac{dy}{c-g(y)}=\frac{dx}{f(x)}\] is what i see it as

OpenStudy (anonymous):

I got until here and then I'm stuck. \[\int\limits_{}^{} 1/1+y*2 dy = \int\limits_{}^{} 1/(1+x)*2 dx\]

OpenStudy (amistre64):

the rest is just integrating tho ...

OpenStudy (amistre64):

your swapping things about looks good ... except that should be a -y^2 over there

OpenStudy (anonymous):

I forgot how to integrate this type of function and I can't find any examples in my books.

OpenStudy (amistre64):

the right side, the xs is just a power rule the left side looks triggy to me

OpenStudy (amistre64):

or since its a difference of squares, it might have to go .... whats the term for splitting the denom?

OpenStudy (amistre64):

partial fraction decomp

OpenStudy (amistre64):

\[1-y^2=(1+y)(1-y)\]

OpenStudy (amistre64):

\[\frac{1}{1-y^2}=\frac{a}{1-y}+\frac{b}{1+y}\] \[1=a(1+y)+b(1-y)\] \[1=a+ay+b-by\] \[1=(a-b)y + (a+b)\] a-b=0 a+b=1 ------ 2a = 1 ; a=1/2 b has to be -1/2 \[\frac{1}{1-y^2}=\frac{1}{2(1-y)}-\frac{1}{2(1+y)}\]

OpenStudy (amistre64):

hmm, i think b = 1/2 not -1/2 :)

OpenStudy (anonymous):

Yes, b= 1/2.

OpenStudy (amistre64):

then after that its back to simple integration

OpenStudy (amistre64):

\[\int(\frac{1}{2(1-y)}+\frac{1}{2(1+y)})dy\]

OpenStudy (amistre64):

partial fraction decomp is what everyone moaned and groaned about in class last semester :)

OpenStudy (anonymous):

I still groan and moan about it till now even though it was last semester's work cause it is related to this semester's topic.

OpenStudy (anonymous):

Hey if y=0 and x=0 then how do I solve it ?

OpenStudy (amistre64):

\[-\frac{1}{2}ln(1-y)+\frac{1}{2}ln(1+y)\] \[\frac{1}{2}[ln(1+y)-ln(1-y)]\] \[\frac{1}{2}ln(\frac{1+y}{1-y})=-\frac{1}{1+x}+C\] \[x=y=0\] \[\frac{1}{2}ln(1)=-1+C\] \[0=-1+C\] \[1=C\]

OpenStudy (amistre64):

once youve integrated; just plug in y and x to determine "C"

OpenStudy (anonymous):

In your fist step, how did you get\[-1/2\ln (1-y)\]. How did you get the (-) sign ? Shouldn't it be a positive value ?

OpenStudy (anonymous):

Can someone please answer me.

OpenStudy (turingtest):

\[u=1-y\implies du=-dy\implies dy=-du\]

OpenStudy (turingtest):

\[\int\frac{dy}{1-y}=-\int\frac{du}u=-\ln u=-\ln|1-y|+C\]

OpenStudy (anonymous):

Ok, thanks.

OpenStudy (amistre64):

i got trapped by the downtime :)